OpenStudy (vortish):

multiply and simplify the expression (9-z)z

OpenStudy (anonymous):

9z-z^2

OpenStudy (vortish):

can you show me how you did that pls

OpenStudy (anonymous):

Distributive property.

OpenStudy (anonymous):

\[ c(b+a)=cb+ca \\ c(b-a)=cb-ca \]

OpenStudy (anonymous):

In this case, \(c=z,b=9,a=z\)

OpenStudy (anonymous):

all you do is multiply th z that is on the outside with what is on the inside